5(5)^(-0.08x)=4

please solve for x

Answer by nerdybill(5404)

5(5)^(-0.08x)=4
If so, start by dividing both sides by 5:
(5)^(-0.08x)=4/5
(5)^(-0.08x)=.8
Take the LOG5 of both sides:
-0.08x = LOG5(.8)
-0.08x = LOG(.8)/LOG(5)
-0.08x = -.0969/.6990
-0.08x = -.1386
x = 1.733